OpenStudy (anonymous):

Find equations of the line tangent to the curve y = (x-1) / (x+1) that are parallel to the line x-2y=2

OpenStudy (anonymous):

First thing to do is find the derivative of the curve. Do you know how to do that?

OpenStudy (anonymous):

second thing to do is to set the result equal to \[\frac{1}{2}\] because that is the slope of your given line

OpenStudy (anonymous):

idk how 2 find teh derivative

OpenStudy (anonymous):

sorry tahts wat confuses me

OpenStudy (anonymous):

ok no problem. i will show you

OpenStudy (anonymous):

Have you heard of the Quotient Rule??

OpenStudy (anonymous):

yes: f'g - fg' /g^2

OpenStudy (anonymous):

Almost

OpenStudy (anonymous):

\[\frac{g(x)(f'(x)) - f(x)(g'(x))}{g(x)^2}\]

OpenStudy (anonymous):

so if you plug in the numbers you get \[\frac{(x+1)(1) - (x-1)(1)}{(x+1)^2}\]

OpenStudy (anonymous):

Do you see how i got that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good

OpenStudy (anonymous):

Distribute the negative sign and you get \[\frac{x + 1 - x + 1}{(x+1)^2}\]

OpenStudy (anonymous):

dont we need to find the derivative of each of the f and g expressions?

OpenStudy (anonymous):

Notice the x's cancel and the ones add up so the derivative is \[\frac{2}{(x+1)^2}\]

OpenStudy (anonymous):

I found the derivatives of f and g when i plugged them into the Quotient rule up at the top. The derivative of (x+1) is 1 and the derivative of (x-1) is also 1.

OpenStudy (anonymous):

The quotient rule says if you have f(x)/g(x).. first put (g(x))^2 in the denominator.. in the numerator keep g(x) the same and multiply by the derivative of f(x) minus f(x) times the derivative of g(x)

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

I apologize, i should of split up the two steps. Do you understand what i did?

OpenStudy (anonymous):

yes..then u simply substituted f and g and their derivatives into the equation

OpenStudy (anonymous):

right and then simplified.

OpenStudy (anonymous):

so now how do we find teh equation of the tangent line

OpenStudy (anonymous):

Ok it says it wants it parallel to the line \[x - 2y = 2\] You need to put this in standard form which is y=mx + b. So can you do that?

OpenStudy (anonymous):

y = 1/2 x +1

OpenStudy (anonymous):

There you go.. Now a line is said to be parallel when two lines have the same slope. So the slope of the tangent line will also be 1/2

OpenStudy (anonymous):

so do we pick any b intercept

OpenStudy (anonymous):

That is the part that im not sure.. I believe you plug in 1/2 into the derivative and that will be your y intercept.

OpenStudy (anonymous):

or set it equal to 1/2.. im not exactly sure.. let me see if i can figure it out

OpenStudy (anonymous):

well wat is the equation of the line that is tangent to the original equation...shudn't we find taht out first

OpenStudy (anonymous):

Just a sec.. i think i figured it out

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Then again maybe not.. Maybe satelite can help you with the last part.. After you find a point on the graph you use the point-slope formula to find the equation.. but im tired and cant think of how to do that.

OpenStudy (anonymous):

its okay..thank you for your help

OpenStudy (anonymous):

but can u explain one last thing

OpenStudy (anonymous):

Whats that?

OpenStudy (anonymous):

how do i find teh equation of the lien tangent to the orig. equation after finding the derivative is equal to 2/ (x+1)^2

OpenStudy (anonymous):

You find the slope which in your case is 1/2 and then you find a y value and then use the point-slope formula.

OpenStudy (anonymous):

is x = 1? since i equated 2/(x+1)^2 = 1/2